OpenStudy (anonymous):

Find an equation of the line with slope negative four thirds and y- intercept 3. Leave equation in slope intercept form. Thanks.

OpenStudy (phi):

replace the m and b in y= mx+b with the numbers they gave you.

OpenStudy (anonymous):

y= 4/3x+3

OpenStudy (phi):

line with slope negative four thirds

OpenStudy (anonymous):

So its's just y= -4/3x+ 3

OpenStudy (phi):

though I like to put parens in y= (-4/3)x+3 because -4/3x looks like \( -\frac{4}{3x}\) and we want \(-\frac{4}{3}x\)